OpenStudy (anonymous):

explain the relationship between the exponential function f(x) = 2^x+1 and it's inverse? urgent.. exam tomorrow D:

OpenStudy (anonymous):

@sourwing @phi please?

OpenStudy (anonymous):

\[ y=2^x+1\\ y=\frac{\ln (x-1)}{\ln (2)} \] They are symmetric with the line y=x

OpenStudy (anonymous):

See the graphs

OpenStudy (anonymous):

i need a detailed explanation D: and I'm not allowed to use my calc for such sections ;_;

OpenStudy (anonymous):

Try to fill in the details. Do you know how to find the inverse of a function?

OpenStudy (anonymous):

f(x)=(2^x) +1 f(x)-1=2^x apply log in base e ln [f(x)-1] =ln(2^x) ln [f(x)-1] =xln2 {1/[f(x)-1]}f'(x)=ln2 f'(x)=ln2 [f(x)-1] f'(x)=ln2[2^x +1-1] f'(x)=ln2(2^x

OpenStudy (anonymous):

f(x)= 2^x+1 let y be the image of x under 'f' so, y= 2^x+1 now, we need to interchange the values of x and y,, in this way: x=2^y+1 or, x-1=2^y or, (root over) x-1 = y hence, f inverse (x) = (root over) x-1

OpenStudy (anonymous):